# Autonomy-of-Experts Models

### v1

@@ Abstract
Mixture-of-Experts (MoE) models mostly use a router to assign tokens to specific expert modules, activating only partial parameters and often outperforming dense models. We argue that the separation between the router’s decision-making and the experts’ execution is a critical yet overlooked issue, leading to suboptimal expert selection and ineffective learning. To address this, we propose Autonomy-of-Experts (AoE), a novel MoE paradigm in which experts autonomously select themselves to process inputs. AoE is based on the insight that an expert is aware of its own capacity to effectively process a token, an awareness reflected in the scale of its internal activations. In AoE, routers are removed; instead, experts pre-compute internal activations for inputs and are ranked based on their activation norms. Only the top-ranking experts proceed with the forward pass, while the others abort. The overhead of pre-computing activations is reduced through a low-rank weight factorization. This self-evaluating-then-partner-comparing approach ensures improved expert selection and effective learning. We pre-train language models having 700M up to 4B parameters, demonstrating that AoE outperforms traditional MoE models with comparable efficiency.
@@ 2 Background: Mixture-of-Experts (MoE)
We focus on studies of sparse MoE models, treating each feed-forward network (FFN) module as an expert. Each FFN, or expert, is expected to possess diverse and distinct abilities, enabling the model to process inputs effectively by activating only the experts with the necessary capabilities, thereby improving efficiency. Some studies (Chen et al., 2024; Lin et al., 2024) on dense MoE do not reduce the parameter activation ratio, which is not the primary concern of this paper. In this paper, when we refer to MoE, we mean sparse MoE.
@@
1:  Input: A hidden state $\mathbf{x}\in\mathbb{R}^{d_{\text{model}}}$, number of experts $n$. Initialize output $\mathbf{h}\in\mathbb{R}^{d_{\text{model}}}$ as a zero vector.
2:  $\mathbf{p}=R(\mathbf{x})$ // $\mathbf{p}\in\mathbb{R}^{n}$
3:  $\mathbf{I}=\texttt{argtopK}(\mathbf{p})$ // $\mathbf{I}\in\mathbb{R}^{K}$
4:  $\mathbf{\hat{p}}=\texttt{Softmax}(\mathbf{p}[\mathbf{I}])$ // $\mathbf{\hat{p}}\in\mathbb{R}^{K}$
5:  for $i=1$ to $n$ do
6:     if $i\in\mathbf{I}$ then
7:        $\mathbf{h}\mathrel{+}=\mathbf{\hat{p}}[i]\cdot E_{i}(\mathbf{x})$
8:     end if
9:  end for
@@
MoE-based LLMs (Jiang et al., 2024; Dai et al., 2024; Team, 2024; Lieber et al., 2024; Sun et al., 2024; Abdin et al., 2024) typically follow the FFN design in the Llama models (Touvron et al., 2023) as an expert module. The $i$-th expert within a specific layer can be formulated as:
@@
Several variants of this auxiliary loss have been proposed (Zuo et al., 2022; Wang et al., 2024a, b; Huang et al., 2024), but they all share the common goal of maintaining a balanced load. Therefore, our discussion focuses on the balancing loss presented above.
Several studies (Roller et al., 2021; Gururangan et al., 2022; Ren et al., 2023; Fan et al., 2020) classify tokens based on prior knowledge—such as domain, language, or hash mapping—and assign them to fixed experts. While they do not use explicit routers, they differ significantly from AoE in many respects. Most importantly, their expert selection is not determined by the experts themselves, leaving the separation between decision-making and execution unaddressed.
@@ 3.1 An Insight: Experts “Know” What They Know
1:  Input: A hidden state $\mathbf{x}\in\mathbb{R}^{d_{\text{model}}}$, number of experts $n$. Initialize output $\mathbf{h}\in\mathbb{R}^{d_{\text{model}}}$, activation cache $\mathbf{C}\in\mathbf{R}^{n\times d_{\text{low}}}$ and $\mathbf{p}\in\mathbb{R}^{n}$ as all zeros.
2:  // In practice, we replace the following loop with a
3:  // single matrix multiplication (see Eq. 4) for efficiency.
4:  for $i=1$ to $n$ do
5:     $\mathbf{C}[i]=\mathbf{x}\mathbf{W}^{i}_{\text{down}}$ // $\mathbf{C}[i]\in\mathbb{R}^{d_{\text{low}}}$
6:  end for
@@
8:  $\mathbf{I}=\texttt{argtopK}(\mathbf{p})$ // $\mathbf{I}\in\mathbb{R}^{K}$
9:  $\mathbf{\hat{p}}=\texttt{Softmax}(\mathbf{p}[\mathbf{I}])$ // $\mathbf{\hat{p}}\in\mathbb{R}^{K}$
10:  for $i=1$ to $n$ do
11:     if $i\in\mathbf{I}$ then
12:        $\mathbf{h}\mathrel{+}=\mathbf{\hat{p}}[i]\cdot\left((\texttt{SiLU}(\mathbf{C}_{i}\mathbf{W}^{i}_{\text{up}})\odot(\mathbf{x}\mathbf{W}^{i}_{p}))\mathbf{W}^{i}_{o}\right)$
13:     end if
14:  end for
@@ 3.2 Autonomy-of-Experts (AoE)
Strategy Model ARC-E PIQA SIQA WINO HELLA MNLI QNLI SST2 AVG.
@@ 4.1.1 General Setup
We conduct a comprehensive evaluation of language models across a range of widely used tasks, including ARC-easy (Clark et al., 2018), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), Winogrande (Sakaguchi et al., 2019), HellaSwag (Zellers et al., 2019), MNLI (Williams et al., 2018), MRPC (Dolan & Brockett, 2005), QNLI (Wang et al., 2019), QQP (Wang et al., 2019), and SST-2 (Socher et al., 2013). The first five tasks are evaluated zero-shot, while the remaining tasks are tested three-shot because models exhibit unstable performance in zero-shot scenarios, with most errors arising from incorrect answer formats. The evaluation code is based on the LM Evaluation Harness (Gao et al., 2024), and accuracy is reported in Table 2.
@@ 4.1.2 Resolving Questions Regarding AoE
Question 1: How does the downstream performance of AoE compare with traditional MoE models?
We evaluated various configurations of AoE (Configs. 5 to 12) and traditional MoE models (Configs. 1 to 4). Every AoE setup outperforms the best-performing MoE configuration in terms of average accuracy across eight tasks. Notably, AoE without any auxiliary loss surpasses traditional MoE models, which enhances the simplicity of training an MoE model. Additionally, AoE exhibits lower training loss, suggesting more efficient training. We elaborate on this in Question 2.
@@
Question 3: Is AoE compatible with other expert-selection strategies?
We also train language models using the Top-$P$ token-choice (Huang et al., 2024) and the Top-$K$ expert-choice strategy (Zhou et al., 2022). Table 3 shows that AoE outperforms traditional MoE models, demonstrating its generality. Details are provided in Appendix B.
Question 4: How is the load balancing of AoE?
@@
Finding 4.1: AoE improves load balancing compared to traditional MoE models, with or without $\mathcal{L}_{\text{aux}}$.
@@
$\mathbf{P}_{i}=\frac{1}{T}\sum_{\mathbf{x}\in\mathcal{B}}\texttt{Softmax}\left(\texttt{L2-Norm}\left(\mathbf{xW}^{i}_{\text{down}}\right)\right)[i].$
$\alpha_{\text{aux}}$ is determined using a validation set comprising 5 billion tokens from (Gokaslan & Cohen, 2019). Experiments indicate that $\alpha_{\text{aux}}=0.01$ is effective for both traditional MoE and AoE models. We adopted this value across all configurations without further hyperparameter tuning.
@@
Finding 4.2: AoE models exhibit stronger confidence in expert selection.
@@
Finding 4.3: Beyond improved load balancing, AoE with $\mathcal{L}_{\text{aux}}$ achieves better downstream performance.
@@
Model ARC-E PIQA SIQA WINO HELLA MNLI QNLI SST2 AVG.
Traditional MoE 53.70 65.40 39.10 51.54 35.80 32.19 49.77 57.00 48.06
AoE 55.98 65.61 39.87 52.57 36.77 35.39 50.05 61.93 49.80
Question 5: Do improvements stem from the factorization of $\mathbf{W}_{g}$? We examined the impact of factorizing the experts’ weight matrix on performance by comparing Configurations 3 and 2. The factorization does not significantly influence performance, as expected in Section 3.2, based on findings that the weights of LLMs are inherently low-rank (Li et al., 2018; Aghajanyan et al., 2021; Hu et al., 2022). Therefore, the improvements observed with AoE are not attributed to the factorization of model weights.
Question 6: Does the improvement of AoE come from involving more parameters in expert selection? We increased the size of the router in MoE to include $n\cdot d_{\text{low}}\cdot d_{\text{model}}$ parameters, ensuring that the number of parameters involved in expert selection remains consistent with that of AoE models. Note that in this setup, traditional MoE models have more activated parameters in total. Comparing Config. 4 and 2, the larger router provides a slight performance benefit. However, every AoE setup still outperforms this configuration. Thus, the improvement in AoE is not primarily due to involving more parameters in expert selection.
Question 7: How aligned are the self-evaluation criteria among experts? In AoE models, each expert independently develops self-evaluation criteria for processing tokens, as reflected in their activation scales. This might raise concerns that some experts could become overly “egoistic,” meaning their internal activations are consistently larger than those of others. For example, one expert might produce activations with norms ranging from 10 to 20, while an “ego” expert produces activations with norms from 20 to 30, leading to biased selections that favor the “ego” expert.
@@
In terms of incremental memory, a smaller $d_{\text{low}}$ requires a larger $d_{\text{wide}}$, thereby increasing the memory consumption of $\mathbf{xW}_{\text{up}}$ to $T\cdot d_{\text{wide}}$, where $T$ is the number of tokens. Conversely, a larger $d_{\text{low}}$ results in a larger activation cache, raising memory usage to $n\cdot T\cdot d_{\text{low}}$. For Configs.  6 to 10, $n\cdot d_{\text{low}}<d_{\text{wide}}$, making the primary memory cost stem from the larger up-projection. In contrast, Config.  11 and 12 satisfy $n\cdot d_{\text{low}}>d_{\text{wide}}$, meaning the increased memory usage is more attributable to the larger activation cache. In terms of throughput reduction, a smaller $d_{\text{low}}$ requires more computational resources for the up-projection, while a larger $d_{\text{low}}$ leads to a higher unused activation cache.
@@ 5 Conclusion
Appendix B Pre-training Language Models Using Alternative Expert-Selection Strategy
We train both small traditional MoE and AoE language models by assigning tokens using the Top-$P$ strategy (Huang et al., 2024). All hyperparameters and dataset details closely follow Section 4.1, except that we replace the Top-$K=2$ strategy with Top-$P=0.6$ following (Wang et al., 2024a). Models utilizing the Top-$P$ strategy require an additional auxiliary loss equivalent to minimizing our introduced $\text{Ent}_{\text{conf}}$ (Eq. 8). This ensures that the model does not learn shortcuts by assigning uniform probabilities to all experts, which would activate too many parameters to achieve lower loss. Following (Huang et al., 2024), we set the weight of this regularization term to $10^{-4}$.
$d_{\text{low}}$ for AoE in these two experiments is 256. Results are shown in Table 3. AoE outperforms traditional MoE models, demonstrating its generality across various expert-selection strategies.
Appendix C Additional Interpretation of AoE’s Advantage

### v2

@@ Abstract
Mixture-of-Experts (MoE) models mostly use a router to assign tokens to specific expert modules, activating only partial parameters and often outperforming dense models. We argue that the separation between the router’s decision-making and the experts’ execution is a critical yet overlooked issue, leading to suboptimal expert selection and ineffective learning. To address this, we propose Autonomy-of-Experts (AoE), a novel MoE paradigm in which experts autonomously select themselves to process inputs. AoE is based on the insight that an expert is aware of its own capacity to effectively process a token, an awareness reflected in the scale of its internal activations. In AoE, routers are removed; instead, experts pre-compute internal activations for inputs and are ranked based on their activation norms. Only the top-ranking experts proceed with the forward pass, while the others abort. The overhead of pre-computing activations is reduced through a low-rank weight factorization. This self-evaluating-then-partner-comparing approach ensures improved expert selection and effective learning. We pre-train language models having 700M up to 4B parameters, demonstrating that AoE outperforms traditional MoE models with comparable efficiency. The code is available at https://github.com/trestad/Autonomy-of-Experts
@@ 2 Background: Mixture-of-Experts (MoE)
We focus on sparse MoE models, treating each feed-forward network (FFN) module as an expert. Each FFN, or expert, is expected to possess diverse and distinct abilities, enabling the model to process inputs effectively by activating only the experts with the necessary capabilities, thereby improving efficiency. Some studies (Chen et al., 2024; Lin et al., 2024) on dense MoE do not reduce the parameter activation ratio, which is not the primary concern of this paper. In this paper, when we refer to MoE, we mean sparse MoE.
@@
0:  A hidden state $\mathbf{x}\in\mathbb{R}^{d_{\text{model}}}$, number of experts $n$.
0:  The layer output $\mathbf{h}\in\mathbb{R}^{d_{\text{model}}}$, initialized as zeros.
1:  $\mathbf{p}=R(\mathbf{x})$ // $\mathbf{p}\in\mathbb{R}^{n}$
2:  $\mathbf{I}=\texttt{argtopK}(\mathbf{p})$ // $\mathbf{I}\in\mathbb{R}^{K}$
3:  $\mathbf{\hat{p}}=\texttt{Softmax}(\mathbf{p}[\mathbf{I}])$ // $\mathbf{\hat{p}}\in\mathbb{R}^{K}$
4:  for $i=1$ to $n$ do
5:     if $i\in\mathbf{I}$ then
6:        $\mathbf{h}\mathrel{+}=\mathbf{\hat{p}}[i]\cdot E_{i}(\mathbf{x})$
7:     end if
8:  end for
@@
MoE-based LLMs (Jiang et al., 2024; Dai et al., 2024; Team, 2024; Lenz et al., 2025; Sun et al., 2024; Abdin et al., 2024) typically follow the FFN design in the Llama models (Touvron et al., 2023) as an expert module. The $i$-th expert within a specific layer can be formulated as:
@@
Several variants of this auxiliary loss are proposed (Zuo et al., 2022; Wang et al., 2024a, b; Huang et al., 2024), sharing the same load-balancing goal. Therefore, our discussion focuses on the balancing loss presented above.
Several studies (Roller et al., 2021; Gururangan et al., 2022; Ren et al., 2023; Fan et al., 2021) classify tokens based on prior knowledge—such as domain, language, or hash mapping—and assign them to fixed experts. While they do not use explicit routers, they differ significantly from AoE in many respects. Most importantly, their expert selection is not determined by the experts themselves, leaving the separation between decision-making and execution unaddressed. Pham et al. (2024) use the norm of expert final outputs as the label for router logits. This method shares the concept with ours, where the activation norm represents expertise; however, it incurs dense activation across all experts and does not address the separation issue we highlighted.
@@ 3.1 An Insight: Experts “Know” What They Know
0:  A hidden state $\mathbf{x}\in\mathbb{R}^{d_{\text{model}}}$, number of experts $n$. Initialize the activation cache $\mathbf{C}\in\mathbf{R}^{n\times d_{\text{low}}}$ and $\mathbf{p}\in\mathbb{R}^{n}$ as all zeros.
0:  The layer output $\mathbf{h}\in\mathbb{R}^{d_{\text{model}}}$, initialized as zeros.
1:  // In practice, we replace the following loop with a
2:  // single matrix multiplication (see Eq. 4) for efficiency.
3:  for $i=1$ to $n$ do
4:     $\mathbf{C}[i]=\mathbf{x}\mathbf{W}^{i}_{\text{down}}$ // $\mathbf{C}[i]\in\mathbb{R}^{d_{\text{low}}}$
5:  end for
@@
7:  $\mathbf{I}=\texttt{argtopK}(\mathbf{p})$ // $\mathbf{I}\in\mathbb{R}^{K}$
8:  $\mathbf{\hat{p}}=\texttt{Softmax}(\mathbf{p}[\mathbf{I}])$ // $\mathbf{\hat{p}}\in\mathbb{R}^{K}$
9:  for $i=1$ to $n$ do
10:     if $i\in\mathbf{I}$ then
11:        $\mathbf{h}\mathrel{+}=\mathbf{\hat{p}}[i]\cdot\left((\texttt{SiLU}(\mathbf{C}_{i}\mathbf{W}^{i}_{\text{up}})\odot(\mathbf{x}\mathbf{W}^{i}_{p}))\mathbf{W}^{i}_{o}\right)$
12:     end if
13:  end for
@@ 4.1.1 General Setup
We conduct a comprehensive evaluation of language models across a range of widely used tasks, including ARC-easy (Clark et al., 2018), PIQA (Bisk et al., 2020), SIQA (Sap et al., 2019), Winogrande (Sakaguchi et al., 2019), HellaSwag (Zellers et al., 2019), MNLI (Williams et al., 2018), MRPC (Dolan & Brockett, 2005), QNLI (Wang et al., 2019), QQP (Wang et al., 2019), and SST-2 (Socher et al., 2013). The first five tasks are evaluated zero-shot, while the remaining tasks are tested three-shot because models exhibit unstable performance in zero-shot scenarios, with most errors arising from incorrect answer formats. The accuracy is reported in Table 2.
@@ 4.1.2 Resolving Questions Regarding AoE
Question 1: How does the downstream performance of AoE compare with traditional MoE models? We evaluated various configurations of AoE (Configs. 5 to 12) and traditional MoE models (Configs. 1 to 4). Every AoE setup outperforms the best-performing MoE setup in terms of average accuracy across eight tasks. Notably, AoE without any auxiliary loss surpasses traditional MoE models, which enhances the simplicity of training an MoE model. Additionally, AoE exhibits lower training loss, suggesting more efficient training. We elaborate on this in Question 2.
@@
Question 3: How is the load balancing of AoE?
@@
Finding 3.1: AoE improves load balancing compared to traditional MoE models, with or without $\mathcal{L}_{\text{aux}}$.
@@
$\mathbf{P}_{i}=\frac{1}{T}\sum_{\mathbf{x}\in\mathcal{B}}\texttt{Softmax}\left(\texttt{L2-Norm}\left(\mathbf{xW}^{i}_{\text{down}}\right)\right)[i],$
and $\alpha_{\text{aux}}$ is determined using a validation set comprising 5 billion tokens from (Gokaslan & Cohen, 2019). Experiments indicate that $\alpha_{\text{aux}}=0.01$ is effective for both traditional MoE and AoE models. We adopted this value across all configurations without further hyperparameter tuning.
@@
Finding 3.2: AoE models exhibit stronger confidence in expert selection.
@@
Finding 3.3: Beyond improved load balancing, AoE with $\mathcal{L}_{\text{aux}}$ achieves better downstream performance.
@@
Strategy Model ARC-E PIQA SIQA WINO HELLA MNLI QNLI SST2 AVG.
@@
Question 4: Do improvements stem from the factorization of $\mathbf{W}_{g}$? We examined the impact of factorizing the experts’ weight matrix on performance by comparing Configurations 3 and 2. The factorization does not significantly influence performance, as expected in Section 3.2, based on findings that the weights of LLMs are inherently low-rank (Li et al., 2018; Aghajanyan et al., 2021; Hu et al., 2022). Therefore, the improvements observed with AoE are not attributed to the factorization of model weights.
Question 5: Does the improvement of AoE come from involving more parameters in expert selection? We increased the size of the router in MoE to include $n\cdot d_{\text{low}}\cdot d_{\text{model}}$ parameters, ensuring that the number of parameters involved in expert selection remains consistent with that of AoE models. Note that in this setup, traditional MoE models have more activated parameters in total. Comparing Config. 4 and 2, the larger router provides a slight performance benefit. However, every AoE setup still outperforms this configuration. Thus, the improvement in AoE is not primarily due to involving more parameters in expert selection.
Question 6: How aligned are the self-evaluation criteria among experts? In AoE models, each expert independently develops self-evaluation criteria for processing tokens, as reflected in their activation scales. This might raise concerns that some experts could become overly “egoistic,” meaning their internal activations are consistently larger than those of others. For example, one expert might produce activations with norms ranging from 10 to 20, while an “ego” expert produces activations with norms from 20 to 30, leading to biased selections that favor the “ego” expert.
@@
Question 7: Is AoE compatible with other expert-selection strategies? We also train language models using the Top-$P$ token-choice (Huang et al., 2024) and the Top-$K$ expert-choice strategy (Zhou et al., 2022).
For Top-$P$ token-choice (Huang et al., 2024), we replace the Top-$K=2$ strategy with Top-$P=0.6$ following (Wang et al., 2024a). Models utilizing the Top-$P$ strategy require an additional auxiliary loss equivalent to minimizing our introduced $\text{Ent}_{\text{conf}}$ (Eq. 8). This ensures that the model does not learn shortcuts by assigning uniform probabilities to all experts, which would activate too many parameters to achieve lower loss. Following (Huang et al., 2024), we set the weight of this regularization term to $10^{-4}$.
Results are shown in Table 3, where $d_{\text{low}}$ in these experiments is 256. AoE outperforms traditional MoE models, demonstrating its generality across various expert-selection strategies.
@@
Model ARC-E PIQA SIQA WINO HELLA MNLI QNLI SST2 AVG.
Traditional MoE 53.70 65.40 39.10 51.54 35.80 32.19 49.77 57.00 48.06
AoE 55.98 65.61 39.87 52.57 36.77 35.39 50.05 61.93 49.80
In terms of incremental memory, a smaller $d_{\text{low}}$ requires a larger $d_{\text{wide}}$, thereby increasing the memory consumption of $\mathbf{xW}_{\text{up}}$ to $T\cdot d_{\text{wide}}$, where $T$ is the number of tokens. Conversely, a larger $d_{\text{low}}$ results in a larger activation cache, raising memory usage to $n\cdot T\cdot d_{\text{low}}$. For Configs.  6 to 10, $n\cdot d_{\text{low}}<d_{\text{wide}}$, making the primary memory cost stem from the larger up-projection. In contrast, Config.  11 and 12 satisfy $n\cdot d_{\text{low}}>d_{\text{wide}}$, meaning the increased memory usage is more attributable to the larger activation cache. In terms of throughput reduction, a smaller $d_{\text{low}}$ requires more computational resources for the up-projection, while a larger $d_{\text{low}}$ leads to a higher unused activation cache. It is worth noting that the efficiency of AoE diminishes as the number of experts increases and as sparsity grows. We are actively working on further optimizing AoE’s efficiency under these conditions.
@@ 5 Conclusion
Acknowledgement
Ang Lv is supported by the Outstanding Innovative Talents Cultivation Funded Programs 2023 of Renmin University of China and CIE-Tencent Doctoral Student Research Incentive Program (HunYuan Large Language Model Special Project). Ruobing Xie is supported by the Young Elite Scientists Sponsorship Program by CAST (2023QNRC001). This work is also supported by the Public Computing Cloud, Renmin University of China and by fund for building world-class universities (disciplines) of Renmin University of China.
Impact Statement
Training large language models can generate content with ethical implications. Many effective techniques can align the preferences and values of LLMs to mitigate these concerns. Beyond this, we believe that our work does not introduce additional societal or ethical issues.
@@ Acknowledgement
Appendix B Additional Interpretation of AoE’s Advantage